# Predicting Polymer Properties Based on Multimodal Multitask Pretraining

### v1

@@ 4.2.2 Adaptability for various downstream settings
Star Substitution 0.431±${}_{\pm\text{0.017}}$ 0.496±${}_{\pm\text{0.031}}$ 0.286±${}_{\pm\text{0.029}}$ 0.390±${}_{\pm\text{0.057}}$ 16.814±${}_{\pm\text{0.867}}$ 0.511±${}_{\pm\text{0.035}}$ 0.087±${}_{\pm\text{0.010}}$ ${}_{\pm\text{0.016}}$
@@
Star Substitution 0.924±${}_{\pm\text{0.006}}$ 0.934±${}_{\pm\text{0.008}}$ 0.925±${}_{\pm\text{0.025}}$ 0.836±${}_{\pm\text{0.053}}$ 0.488±${}_{\pm\text{0.072}}$ 0.779±${}_{\pm\text{0.052}}$ 0.864±${}_{\pm\text{0.036}}$ 0.961±${}_{\pm\text{0.018}}$
4.3 Ablation Studies
4.3.1 Star Substitution Strategy
In particular, the performance of MMPolymer degrades when using the "Star Remove" strategy. We attribute this phenomenon to the lack of polymerization site information. As mentioned in Sec. 3.2.1, the P-SMILES string is formed by combining the SMILES string of the corresponding monomer with two "*" symbols, where "*" symbols are used to indicate the polymerization sites. Therefore, removing "*" symbols means using the monomer’s 3D conformation rather than the repeating unit’s 3D conformation. In this case, it’s difficult to capture the structural features of the whole polymer 3D conformation, thus leading to worse performance. This phenomenon also further indicates the specificity of polymers and the need for tailored methods rather than directly transferring methods from other scientific domains (e.g., protein domain or molecule domain).
Table 5: The performance of MMPolymer under different pretraining settings, where "1D-pre" refers to the masking prediction task on the 1D representation network, "3D-pre" refers to the coordinate denoising task on the 3D representation network, and "Contr" refers to the cross-modal alignment task.
Metric 1D-pre 3D-pre Contr Egc Egb Eea Ei Xc EPS Nc Eat
@@ 4.3.1 Star Substitution Strategy
4.3.2 Multimodal Multitask Pretraining Paradigm
@@
Specifically, we observe that incorporating 3D structural information greatly enhances the performance of MMPolymer, especially when combining both 1D sequential and 3D structural information, showcasing the critical role of 3D structural information in polymer property prediction. Moreover, the utilization of contrastive learning further enhances the performance of MMPolymer. This suggests that aligning learned representations across different modalities contributes to better multimodal integration, thus improving the predictive performance of polymer properties. Notably, we aim to leverage complementary 1D and 3D information for polymer property prediction, rather than matching the two modalities. In this case, purely applying the contrastive learning task is insufficient and leads to unsatisfactory performance.
@@
Besides, we observe that the performance of MMPolymer with only 1D pretraining surpasses that with 3D pretraining on most datasets. We attribute this phenomenon to the fact that we approximate the whole polymer 3D conformation with the 3D conformation of its corresponding repeating unit. Although through our "Star Substitution" strategy, as analyzed in Sec. 4.3.1, the 3D conformation of the repeating unit can reflect the structural features of the whole polymer 3D conformation to some extent, it is still not accurate enough to describe all natures. This observation highlights the current issue of insufficient high-precision polymer 3D structure data and indicates that pure 3D pretraining methods still face significant challenges in the polymer domain, further emphasizing the benefits of our multimodal multitask pretraining paradigm.
5 Conclusions
@@
In this work, we present MMPolymer, a multimodal multitask pretraining framework, to achieve accurate polymer property prediction. By effectively combining polymer 1D sequential and 3D structural information through our multimodal multitask pretraining paradigm, MMPolymer can fully capture diverse aspects of polymer data, creating a robust and promising model for the downstream property prediction tasks. Besides, through our "Star Substitution" strategy, the 3D structural information can be extracted effectively, overcoming the scarcity of polymer 3D data to some extent. The extensive experiments demonstrate that MMPolymer achieves state-of-the-art performance on various polymer property prediction tasks, significantly outperforming existing methods. Even if a single modality (P-SMILES string or 3D conformation) is utilized during fine-tuning, the pretrained MMPolymer can still surpass existing methods, showcasing its exceptional capability in polymer feature extraction and utilization. Moreover, comprehensive ablation studies are also conducted, providing valuable insights into the rationality of our method. As polymer research continues to advance, our proposed MMPolymer provides a valuable strategy for accurate polymer property prediction, paving the way for further developments in this field. For future work, we will continue to explore the intrinsic relationship between polymer structures and their properties, achieving better modeling of polymer structures to benefit its many downstream applications.

### v2

@@ 4.1.4. Model architecture and hyperparameter settings
Star Substitution (Ours) 0.431±${}_{\pm\text{0.017}}$ 0.496±${}_{\pm\text{0.031}}$ 0.286±${}_{\pm\text{0.029}}$ 0.390±${}_{\pm\text{0.057}}$ 16.814±${}_{\pm\text{0.867}}$ 0.511±${}_{\pm\text{0.035}}$ 0.087±${}_{\pm\text{0.010}}$ ${}_{\pm\text{0.016}}$
@@
Star Substitution (Ours) 0.924±${}_{\pm\text{0.006}}$ 0.934±${}_{\pm\text{0.008}}$ 0.925±${}_{\pm\text{0.025}}$ 0.836±${}_{\pm\text{0.053}}$ 0.488±${}_{\pm\text{0.072}}$ 0.779±${}_{\pm\text{0.052}}$ 0.864±${}_{\pm\text{0.036}}$ 0.961±${}_{\pm\text{0.018}}$
Table 5. The performance of MMPolymer under different pretraining settings, where "1D pre" refers to the masking prediction task on the 1D representation network, "3D pre" refers to the coordinate denoising task on the 3D representation network, and "Contrast" refers to the cross-modal alignment task.
Metric 1D pre 3D pre Contrast Egc Egb Eea Ei Xc EPS Nc Eat
@@
4.2. Comparisons
4.2.1. Demonstration of effectiveness
Table 2 presents the performance of different methods on the eight polymer property datasets, which are evaluated by RMSE and $R^{2}$. In general, our proposed method, MMPolymer, stands out by achieving state-of-the-art performance on seven datasets and comparable performance on the remaining Eat dataset, underscoring its effectiveness on polymer property prediction tasks. Notably, several previous works, such as Transpolymer (Xu et al., 2023), have excluded the Eat dataset from their original paper because it fails to capture the complete polymer structure [^3] Eat is a localized property, which depends on the local atomic environment. and therefore leads to lower reliability compared to other property datasets. Besides, even if considering the Eat dataset, the superiority of our MMPolymer is still obvious.
Moreover, comparing the four representative molecular pretraining methods reveals some insightful trends. On the one hand, the 3D molecular pretraining method Uni-Mol consistently outperforms the 1D molecular pretraining method ChemBERTa across all eight polymer property datasets, emphasizing the crucial role of 3D structural information in enhancing predictive performance. On the other hand, all molecular pretraining methods exhibit inferior performance when compared with our MMPolymer. These observations indicate the inherent limitations of directly applying molecular pretraining methods to polymer-specific tasks, emphasizing the significance of tailored methods like our MMPolymer for achieving accurate polymer property prediction.
As shown in Figure 4, we conduct t-SNE visualization (Van der Maaten and Hinton, 2008) based on the polymer representations learned by our MMPolymer and corresponding ground truth for each polymer property dataset. The t-SNE visualization illustrates that the chemical space defined by the representations accurately captures the differences in the ground truth for each data point, as reflected by the color gradations. These colors are determined by the corresponding ground truth, and the spread of colors across the datasets indicates a diverse range of values. This variation confirms that the representations learned by our MMPolymer are sensitive to the underlying ground truth differences among data points. The distinct clusters, especially in datasets like Egb, Eea, and Eat, where the color intensity marks out separate regions, serve as a testament to the fidelity with which the representations map out the actual property landscape of the dataset. The linear color variation further emphasizes that the representations are not only capturing distinct property values but are also able to reflect a continuum in the property and chemical space, showing the superiority of our MMPolymer.
4.2.2. Adaptability for various downstream settings
4.3. Ablation Studies
4.3.1. Star Substitution Strategy
In particular, the performance of MMPolymer degrades when using the "Star Remove" strategy. We attribute this phenomenon to the lack of polymerization site information. As mentioned in Sec. 3.2.1, the P-SMILES string is formed by combining the SMILES string of the corresponding monomer with two "*" symbols, where "*" symbols are used to indicate the polymerization sites. Therefore, removing "*" symbols means using the monomer’s 3D conformation rather than the repeating unit’s 3D conformation. In this case, it’s difficult to capture the structural features of the whole polymer 3D conformation, thus leading to worse performance. This phenomenon also further indicates the specificity of polymers and the need for tailored methods rather than directly transferring methods from other domains (e.g., protein domain or molecule domain).
4.3.2. Multimodal Multitask Pretraining Paradigm
@@
Specifically, we observe that incorporating 3D structural information greatly enhances the performance of MMPolymer, especially when combining both 1D sequential and 3D structural information, showcasing the critical role of 3D structural information in polymer property prediction. Moreover, the utilization of contrastive learning further enhances the performance of MMPolymer. This suggests that aligning learned representations across different modalities contributes to better multimodal integration, thus improving the predictive performance of polymer properties. Notably, we aim to leverage complementary 1D and 3D information for polymer property prediction, rather than matching the two modalities. In this case, purely applying the contrastive learning task is insufficient and leads to unsatisfactory performance, as shown in Table 5.
@@
Besides, we can observe that the performance of MMPolymer with only 1D pretraining surpasses that with 3D pretraining on most datasets. We attribute this phenomenon to the fact that we approximate the whole polymer 3D conformation with the 3D conformation of its corresponding repeating unit. Although through our "Star Substitution" strategy, as analyzed in Sec. 4.3.1, the 3D conformation of the repeating unit can reflect the structural features of the whole polymer 3D conformation to some extent, it is still not accurate enough to describe all natures. This observation highlights the current issue of insufficient high-precision polymer 3D structure data and indicates that pure 3D pretraining methods still face significant challenges in the polymer domain, further emphasizing the benefits of our multimodal multitask pretraining paradigm.
5. Conclusions
@@
In this work, we present MMPolymer, a multimodal multitask pretraining framework, to achieve accurate polymer property prediction [^4] Code is available at https://github.com/FanmengWang/MMPolymer. By effectively combining polymer 1D sequential and 3D structural information through our multimodal multitask pretraining paradigm, MMPolymer can fully capture diverse aspects of polymer data, creating a robust and promising model for downstream polymer property prediction tasks. Besides, through our "Star Substitution" strategy, the 3D structural information can be extracted effectively, overcoming the scarcity of polymer 3D data to some extent. The extensive experiments consistently demonstrate that MMPolymer achieves state-of-the-art performance on various polymer property prediction tasks, significantly outperforming existing polymer property prediction methods. Even if only single-modal information (either polymer 1D sequential or 3D structural information) is utilized in the fine-tuning phase, the pretrained MMPolymer can still surpass existing polymer property prediction methods, showcasing its exceptional capability in polymer feature extraction and utilization. Moreover, comprehensive ablation studies are also conducted to confirm the rationality and effectiveness of our proposed method. For future work, we will continue to explore the intrinsic relationship between polymer structures and their properties, achieving better modeling of polymer structures to benefit its many downstream applications.
Acknowledgements.
This work was supported in part by the National Natural Science Foundation of China (62106271, 92270110), the Fundamental Research Funds for the Central Universities, and the Research Funds of Renmin University of China. Dr. Hongteng Xu thanks the support from the Beijing Key Laboratory of Big Data Management and Analysis Methods, the Intelligent Social Governance Platform, Major Innovation & Planning Interdisciplinary Platform for the “Double-First Class” Initiative.